# RBI-ThPhys-2024-15

### v1

@@ VII.2 BM2 and BM3
Figure 29: The di-Higgs mass distributions for BM2 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta.
Figure 30: The di-Higgs mass distributions for BM3 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta.
The $m_{hh}$ distributions for BM2 and BM3 are shown in Figs. 29 and 30, respectively. In general, the effect of the interference in these BMs is visibly smaller than for BM1, especially after the experimental smearing is applied. However, comparing the ratios in the figures we observe that the relative differences away from the resonant peak are actually quite similar to BM1, but as the relative contribution of the non-resonant component compared to the $\mathrm{S_{H}}$ contribution is smaller the effect is less noticeable.
@@
Since it is possible that the experimental collaborations could improve the $m_{hh}$ resolution in the future, we investigated if such an improvement would significantly change the conclusion drawn in this case. To this end, we artificially improved the $m_{hh}$ resolution by a factor of two and display the resulting $m_{hh}$ distribution in Fig. 31. In this case, we observe a slightly larger difference between the full spectrum and the no-interference scenario in this case.
Figure 31: The di-Higgs mass distributions for BM3 after experimental smearing while artificially improving the $m_{hh}$ resolution by a factor of two compared to Fig. 30. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta.
@@ VII.3 BM4 and BM5
Figs. 32 and 33 show the $m_{hh}$ distributions in BM4 and BM5, respectively. In both cases there is a significant difference between the $\mathrm{S_{H}}$-only assumption and both the full model and no-interference scenarios. BM4 also shows a sizable effect due to the interference. In contrast, the impact of the interference in BM5 is quite small. However, the non-resonant contribution is even more significant in this case and it clearly has to be taken into account by future experimental searches.
Figure 32: The di-Higgs mass distributions for BM4 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta.
Figure 33: The di-Higgs mass distributions for BM5 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta.
@@ VII.4 BM6 and BM7
The $m_{hh}$ distributions are shown in Figs. 34 and 35. Both BMs predict a more sizable contribution to the total di-Higgs spectrum from the non-resonant contribution compared to $\mathrm{S_{H}}$. However, we note that, as the resonant part peaks at higher mass, it may still drive the sensitivity for the experimental searches, as the backgrounds are probably also small for large $m_{hh}$. Once again, these distributions underline the need to take the non-resonant contributions into account in the experimental searches.
Figure 34: The di-Higgs mass distributions for BM6 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta.
Figure 35: The di-Higgs mass distributions for BM7 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta.
@@ VII.5 BM8
The final BM, BM8, is defined as the point which has the largest deviation of $\lambda_{hhh}$ from the SM value. This main purpose of this BM is to investigate whether or not assuming SM-like couplings for the non-resonant contribution is a reasonable approximation. This may have practical implications for experimental searches, i.e. fixing these couplings to SM values reduces the dimensionality of the parameter space that needs to be scanned. The largest deviation for $\lambda_{hhh}$ is found for $m_{H}=260\,{\,\rm GeV}$, $\sin\alpha=0.24$, and $\tan\beta=3.5$ which yields a value of $\kappa_{\lambda_{hhh}}=0.87$.
@@
The $m_{hh}$ distributions are shown in Fig. 36. We observe no major differences between the no-interference and SM+$\mathrm{S_{H}}$ distributions, implying that assuming SM couplings for the $h$ is a reasonable assumption in this case. The non-resonant contribution to the spectrum is less noticeable in these plots as the resonant $\mathrm{S_{H}}$ peak is large. To study in more detail how the values of the $h$ couplings influence the non-resonant contribution, we remove the $\mathrm{S_{H}}$, $-\mathrm{S_{H}}{\text{-}}\Box$, and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ contributions to the spectrum, as shown in Fig. 37. The non-resonant spectrum for BM8 shown in blue is compared to the SM shown in red. We see only small differences between the two distributions, mainly impacting the low $m_{hh}$ bins where the event yield is low. The differences in the distributions are somewhat smaller than one might expect given that $\kappa_{\lambda_{hhh}}$ deviates from the SM value by a non-negligible amount. This is due to a small deviation in the value of $\kappa_{q}^{h}$, which is equal to 0.97 in this BM. This causes an accidental cancellation between the $\Box$ and $\mathrm{S_{h}}$ components, which decrease in size for couplings $<1$, and the $-\mathrm{S_{h}}{\text{-}}\Box$ component which increases due to the destructive nature of the interference. The scenario where only $\kappa_{\lambda_{hhh}}$ deviates from SM and $\kappa_{q}^{h}=1$ is shown in green in the Figure for comparison. Even in this scenario the deviation in the non-resonant yield is only $\approx 10\%$ so it is unlikely that the LHC experiments will be sensitive to the deviations in $\kappa_{\lambda_{hhh}}$ in this model [^17] We note that BM8 has since been excluded by the new combination presented by CMS in Ref. CMS:2024phk that is not yet included in HiggsBounds. However, as the purpose of this benchmark is to study the most extreme cases where $\kappa_{\lambda_{hhh}}$ deviates from the SM, we still consider this a useful scenario since the new bounds from CMS restrict $\kappa_{\lambda_{hhh}}$ to be even closer to unity, and thus the conclusion drawn from this BM are still valid. We thank H. Bahl for useful discussions regarding this point..
@@
Figure 36: The di-Higgs mass distributions for BM8 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta.
@@ VIII Summary
We have presented the latest experimental and theoretical limits on the parameter space of the real singlet extension of the SM Higgs sector. The tightest bounds on the model at high $m_{H}$ come from the measurements of the $W$ boson mass, whereas the bounds at low $\sin\alpha$ are driven by direct searches for $H\rightarrow VV/hh$ at the LHC.
We have investigated in detail interference effects between resonant and non-resonant contributions, where the former originates from the production of a heavy additional Higgs boson which subsequently decays into two lighter SM-like Higgs bosons ($H\rightarrow hh$). We have studied various kinematic observables, as e.g. the commonly investigated di-Higgs production spectrum ($m_{hh}$), as well as a number of other kinematic variables that are relevant for the respective experimental searches. In particular we find $p_{T}^{h}$ to have a comparable sensitivity to interference effects.
We have identified regions of the parameter space where interference effects between the non-resonant and resonant diagrams are sizable. In order to account for NLO kinematics as well as interference effects, we have proposed a prescription that includes both, where we find good agreement with the correct NLO description when available. Using our methodology, we have defined benchmark scenarios exhibiting interesting features of the model. For several benchmarks scenarios, we find that the contribution of the non-resonant di-Higgs diagrams is significant and in some cases even dwarfs the resonant component. We have also observed that interference effects have a non-negligible effect on both the cross-sections and differential distributions. We strongly encourage the experimental collaborations to take these effects into account in future searches for additional Higgs bosons. Furthermore, the correct physical prediction of the total width needs to be used in realistic simulation as the width can crucially impact the rates as well as the relevance of the interference term. Finally, we have presented a publicly available tool, HHReweighter reweight_repo , for performing matrix-element reweighting that allows such interference effects to be taken into account in a computationally efficient way.

### v2

@@ VII.2 BM2 and BM3
Figure 29: The di-Higgs mass distributions for BM2 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta.
Figure 30: The di-Higgs mass distributions for BM3 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta.
The $m_{hh}$ distributions for BM2 and BM3 are shown in Figures 29 and 30, respectively. In general, the effect of the interference in these BMs is visibly smaller than for BM1, especially after the experimental smearing is applied. However, comparing the ratios in the figures we observe that the relative differences away from the resonant peak are actually quite similar to BM1, but as the relative contribution of the non-resonant component compared to the $\mathrm{S_{H}}$ contribution is smaller the effect is less noticeable.
@@
Since it is possible that the experimental collaborations could improve the $m_{hh}$ resolution in the future, we investigated if such an improvement would significantly change the conclusion drawn in this case. To this end, we artificially improved the $m_{hh}$ resolution by a factor of two and display the resulting $m_{hh}$ distribution in Figure 31. In this case, we observe a slightly larger difference between the full spectrum and the no-interference scenario.
Figure 31: The di-Higgs mass distributions for BM3 after experimental smearing while artificially improving the $m_{hh}$ resolution by a factor of two compared to Figure 30. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta.
@@ VII.3 BM4 and BM5
Figures 32 and 33 show the $m_{hh}$ distributions in BM4 and BM5, respectively. In both cases there is a significant difference between the $\mathrm{S_{H}}$-only assumption and both the full model and no-interference scenarios. BM4 also shows a sizeable effect due to the interference. In contrast, the impact of the interference in BM5 is quite small. However, the non-resonant contribution is even more significant in this case and it clearly has to be taken into account by future experimental searches.
Figure 32: The di-Higgs mass distributions for BM4 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta.
Figure 33: The di-Higgs mass distributions for BM5 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta.
@@ VII.4 BM6 and BM7
The $m_{hh}$ distributions are shown in Figures 34 and 35. Both BMs predict a more sizeable contribution to the total di-Higgs spectrum from the non-resonant contribution compared to $\mathrm{S_{H}}$. However, we note that, as the resonant part peaks at higher mass, it may still drive the sensitivity for the experimental searches, as the backgrounds are probably also small for large $m_{hh}$. Once again, these distributions underline the need to take the non-resonant contributions into account in the experimental searches.
Figure 34: The di-Higgs mass distributions for BM6 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta.
Figure 35: The di-Higgs mass distributions for BM7 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta.
@@ VII.5 BM8
BM8 is defined as the point which has the largest deviation of $\lambda_{hhh}$ from the SM value at LO. This main purpose of this BM is to investigate whether or not assuming SM-like couplings for the non-resonant contribution is a reasonable approximation. This may have practical implications for experimental searches, i.e. fixing these couplings to SM values reduces the dimensionality of the parameter space that needs to be scanned. The largest deviation for $\lambda_{hhh}$ is found for $m_{H}=260\,{\,\rm GeV}$, $\sin\alpha=0.24$, and $\tan\beta=3.5$ which yields a value of $\kappa_{\lambda_{hhh}}=0.87$.
@@
The $m_{hh}$ distributions are shown in Figure 36. We observe no major differences between the no-interference and SM+$\mathrm{S_{H}}$ distributions, implying that assuming SM couplings for the $h$ is a reasonable assumption in this case. The non-resonant contribution to the spectrum is less noticeable in these plots as the resonant $\mathrm{S_{H}}$ peak is large. To study in more detail how the values of the $h$ couplings influence the non-resonant contribution, we remove the $\mathrm{S_{H}}$, $-\mathrm{S_{H}}{\text{-}}\Box$, and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ contributions to the spectrum, as shown in Figure 37. The non-resonant spectrum for BM8 shown in blue is compared to the SM shown in red. We see only small differences between the two distributions, mainly impacting the low $m_{hh}$ bins where the event yield is low. The differences in the distributions are somewhat smaller than one might expect given that $\kappa_{\lambda_{hhh}}$ deviates from the SM value by a non-negligible amount. This is due to a small deviation in the value of $\kappa_{q}^{h}$, which is equal to 0.97 in this BM. This causes an accidental cancellation between the $\Box$ and $\mathrm{S_{h}}$ components, which decrease in size for couplings $<1$, and the $-\mathrm{S_{h}}{\text{-}}\Box$ component which increases due to the destructive nature of the interference. The scenario where only $\kappa_{\lambda_{hhh}}$ deviates from the SM and $\kappa_{q}^{h}=1$ is shown in green in the Figure for comparison. Even in this scenario the deviation in the non-resonant yield is only $\approx 10\%$ so it is unlikely that the LHC experiments will be sensitive to the deviations in $\kappa_{\lambda_{hhh}}$ in this model [^18] We note that BM8 has since been excluded by the new combination presented by CMS in Ref. CMS:2024phk that is not yet included in HiggsTools. However, as the purpose of this benchmark is to study the most extreme cases where $\kappa_{\lambda_{hhh}}$ deviates from the SM, we still consider this a useful scenario since the new bounds from CMS restrict $\kappa_{\lambda_{hhh}}$ to be even closer to unity, and thus the conclusions drawn from this BM are still valid. We thank H. Bahl for useful discussions regarding this point. .
@@
Figure 36: The di-Higgs mass distributions for BM8 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta.
@@
VII.6 BM9
The final BM, BM9, is defined as the point which has the largest value of $\frac{\Gamma_{H}}{m_{H}}$, which is found for $m_{H}=800\,{\,\rm GeV}$, $\sin\alpha=0.16$, and $\tan\beta=1.0$. The value of $\frac{\Gamma_{H}}{m_{H}}$ is 1.2%. The $m_{hh}$ distributions are shown in Figures 38.
The main purpose of this BM is to investigate whether finite width effects may invalidate the procedure followed by several experimental searches whereby a narrow width is assumed for the generated $H$ signal. To investigate this question, we display the $m_{hh}$ distributions for the $\mathrm{S_{H}}$ contribution in Figure 39. The figure compares the $m_{hh}$ distributions predicted for samples generated with the correct width to samples generated assuming a narrow-width of $20\,{\rm MeV}$. For the latter, the sample yield is scaled by $(20\,{\rm MeV})/\Gamma_{H}$ to account for the different branching fractions. The distributions are shown for our nominal assumption about the experimental resolution (left plot) and a more optimistic experimental resolution, which is improved by a factor of two (right plot).
The $m_{hh}$ distributions obtained for the nominal smearing are in quite good agreement. The difference between the two distributions is only around 7% at the maximum of the peak, and thus the experiments would probably not be able to resolve the differences. The optimistic scenario is intended to investigate the impact of potential future improvements to the experimental resolution if the experimental collaborations are able to improve the $m_{hh}$ resolution in the future and/or if a channel with a better resolution is used instead (e.g. $bb\gamma\gamma$). As expected, when the resolution is bettered the experiments become more sensitive to the width effects, and the difference between the $m_{hh}$ distributions can be as large as 14% at the maximum of the peak. While this may lead to small modifications to the measured experimental limits, it would probably not significantly alter the overall conclusions of the searches and thus we conclude that the narrow-width assumption for generation of the $\mathrm{S_{H}}$ contributions is a reasonable approximation for the singlet model. However, we stress that this conclusion is specific to the singlet model since other beyond the SM scenarios can still accommodate more sizeable widths $\mathcal{O}(10\%)$, and these scenarios will require a proper treatment of finite width effects. It is also important to stress that, while the finite width effects for the $\mathrm{S_{H}}$-only process can be accounted for via the branching fraction rescaling, this is not possible for the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms. As demonstrated previously in Figure 9, the width modifies these distributions in a non-trivial way (e.g. differentially in $m_{hh}$), and thus it is not possible to account for this by the application of a constant scale factor as for the $\mathrm{S_{H}}$ contribution.
Figure 38: The di-Higgs mass distributions for BM9 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta.
Figure 39: The di-Higgs mass distributions for BM9 after experimental smearing are displayed for the $s$-channel only ($\mathrm{S_{H}}$) contributions. The blue lines show distributions obtained using the correct width ($\Gamma_{H}=9.8\,{\,\rm GeV}$), while the red lines display the distributions obtained using the NWA with $\Gamma_{H}=20\,{\rm MeV}$ and rescaling the the branching fractions by $(20\,{\rm MeV})/\Gamma_{H}$ as described in the text. The left plot shows the distributions for the nominal assumed experimental resolution while the right plot shows the distributions after artificially improving the $m_{hh}$ resolution by a factor of two.
@@ VIII Summary
We have presented the latest experimental and theoretical limits on the parameter space of the real singlet extension of the SM Higgs sector. The tightest bounds on the model at high $m_{H}$ come from the measurements of the $W$-boson mass, whereas the bounds at low $\sin\alpha$ are driven by direct searches for $H\rightarrow VV/hh$ at the LHC.
We have investigated in detail interference effects between resonant and non-resonant contributions, where the former originates from the production of a heavy additional Higgs boson which subsequently decays into two lighter SM-like Higgs bosons ($H\rightarrow hh$). We have studied various kinematic observables, such as the commonly investigated di-Higgs invariant mass ($m_{hh}$), as well as a number of other variables that are relevant for the respective experimental searches. In particular, we find that the transverse momenta of the two SM-like Higgs bosons have a comparable sensitivity to interference effects.
We have identified regions of the parameter space where interference effects between the non-resonant and resonant diagrams are sizeable. In order to account for NLO kinematics as well as interference effects, we have proposed a prescription that includes both, where we find good agreement with the correct NLO description when available. Using our methodology, we have defined benchmark scenarios exhibiting interesting features of the model. For several benchmarks scenarios, we find that the contribution of the non-resonant di-Higgs diagrams is significant and in some cases even dwarfs the resonant component. We have also observed that interference effects have a non-negligible effect on both the cross-sections and differential distributions. We strongly encourage the experimental collaborations to take these effects into account in future searches for additional Higgs bosons. Furthermore, the correct physical prediction of the total width needs to be used in realistic simulation as the width can crucially impact the rates as well as the relevance of the interference term. Finally, we have presented a publicly available tool, HHReweighter reweight_repo , for performing matrix-element reweighting that allows such interference effects to be taken into account in a computationally efficient way.
Appendix A Interference evaluation in an additional observable
In Figure A.1, we display the differential distribution in $\Delta\phi_{hh}$ for both parameter points discussed in Section V.1. The similarity of the blue line (coherent sum of all contributions to $pp\to hh$) and the green line (incoherent sum of the contributions via $H$ and without $H$, i.e. without the interference term) indicates that the interference does not contribute significantly to this variable.
Appendix B Additional validation plots of the reweighting method
The reweighting method described in Section VI is validated for all benchmark points given in Section VII. Figures B.1B.3 show the reweighted $m_{hh}$ distribution (blue) compared to the $m_{hh}$ distribution obtained from a MC sample generated directly for the target model parameters (black).
Figure B.3: The reweighting method is validating by comparing reweighted di-Higgs mass distribution (blue) to the distribution obtained by directly simulating the di-Higgs events for a set of parameter points (black). The distribution is shown for BM9.